2/3 (11x - 1/6) - 3/4 = 1/4 I need to cancel fractions.
Answer by checkley77(12844)

2/3(11x-1/6)-3/4=1/4
22x/3-2/18-3/4=1/4
22x/3=1/4+1/9+3/4
22x/3=(1*9+1*4+3*9)/(4*9)
22x/3=(9+4+27)/36
22x/3=40/36
22x*36=3*40
792x=120
x=120/792
x=5/33 ans.